# Non-Variational Quantum Random Access Optimization with Alternating Operator Ansatz

### v1

@@ Abstract
Solving hard optimization problems is one of the most promising application domains for quantum computers due to the ubiquity of such problems in industry and the availability of broadly applicable quantum speedups. However, the ability of near-term quantum computers to tackle industrial-scale optimization problems is limited by their size and the overheads of quantum error correction. Quantum Random Access Optimization (QRAO) has been proposed to reduce the space requirements of quantum optimization. However, to date QRAO has only been implemented using variational algorithms, which suffer from the need to train instance-specific variational parameters, making them difficult to scale. We propose and benchmark a non-variational approach to QRAO based on the Quantum Alternating Operator Ansatz (QAOA) for the MaxCut problem. We show that instance-independent “fixed” parameters achieve good performance, removing the need for variational parameter optimization. Additionally, we evaluate different design choices, such as various mixers and initial states, as well as QAOA operator implementations when customizing for QRAO, and identify a strategy that performs well in practice. Our results pave the way for the practical execution of QRAO on early fault-tolerant quantum computers.
@@ I Introduction
Combinatorial optimization (CO) is widely considered one of the most promising application domains for quantum computers due to the ubiquity of hard optimization problems in industry and the availability of broadly applicable quantum algorithmic speedups [1, 2, 3, 4, 5, 6]. An important limitation of near-term quantum computers is the small number of computational qubits, which is due to the challenges in scaling up the quantum devices and the space overhead of error correction.
Quantum Random Access Optimization (QRAO) [7, 8] is a quantum-relaxation based optimization method that reduces the space requirements of quantum optimization algorithms by encoding more than one variable per qubit. It utilizes the quantum superposition property, namely, a qubit can be in a superposition of 0 and 1 states, and turns the noise into advantage. That is, while a qubit is necessary to encode a binary variable without error by Holevo bound, it can encode more variables if error is tolerated. Quantum Random Access Codes (QRACs) [9, 10] are the family of such encodings; $(m,N,p)$-QRAC is an encoding of $m$ binary variables on $N$ qubits so that any 1 out of $m$ can be extracted with probability $p>1/2$. It is known that there exist $(4^{N}-1,N,p)$-QRACs but not $(4^{N},N,p)$-QRACs [11], while optimal construction of QRACs is only known for $N=1,2$ [11, 12, 13]. QRACs were first used for CO in [7], extended theoretically in [8], and since then there have been various research revealing QRAO’s advantages; it can scale up the CO instances solvable on near-term quantum devices [14, 15], be robust to noise [16], applicable to constrained CO [17], and its recursive variant can achieve better solutions than other state-of-the-art quantum approaches [18].
@@ II-B Quantum Random Access Optimization (QRAO) formulation
QRAO is based on the idea of encoding multiple binary variables in a qubit. In this paper, we focus on encoding 3 binary variables into a qubit using the $(3,1,1/2+1/(2\sqrt{3})$-QRAC that encodes the bit string $b_{1}b_{2}b_{3}\in\{0,1\}^{3}$ into the quantum state (in the density matrix representation [30])
@@
The main idea of QRAO for solving a $3N$-variable CO instance is to construct a cost Hamiltonian $\tilde{H}_{C}$ of an $N$-qubit system so that its ground state closely resembles $\rho_{b^{*}}\equiv\rho_{b_{1}^{*}b_{2}^{*}b_{3}^{*}}\otimes\cdots\otimes\rho_{b_{3N-2}^{*}b_{3N-1}^{*}b_{3N}^{*}}$, where $b^{*}$ is the optimal solution of the CO instance. The cost Hamiltonian of QRAO for MaxCut $G(V,E)$, here denoted by $\tilde{H}_{C}$, is easily generalized from that of QAOA in Eq. (2) as
@@
However, because the ground state $\rho$ of $\tilde{H}_{C}$ is likely to be a (pure) entangled state found on larger subspace than $\rho_{b}$, $\mbox{Tr}\left(\rho\tilde{H}_{C}\right)\geq\mbox{Tr}\left(\rho_{b}\tilde{H}_{C}\right)$ holds and some rounding procedure, such as the Pauli rounding, must be performed to find the closest $\rho_{b}$ to $\rho$. For this reason, $\tilde{H}_{C}$ is also called relaxed Hamiltonian. Other rounding methods are the magic rounding [7], that randomly projects each qubit of $\rho$ into one-out-of-eight state of Eq. (8), and the tree rounding [18].
@@ III-B Parameter Setting Protocol
Figure 5: Performance of QAOA-for-QRAO with fixed parameters. A We present the ratio of the performance with fixed parameters to the performance with optimized parameters. The fixed parameters performance is comparable to the ones with optimized parameters across different values of $p$ in both $\alpha_{r}$ (blue line) and $\alpha_{c}$ (orange line). B We show the average approximation ratios $\alpha_{r}$ and $\alpha_{c}$ for 120 random instances, with $N$ varying from 10 to 16 and $p$ up to 6. The performance improves as $p$ increases, indicating that the fixed parameters are of high quality across different values of $p$. C We present the average approximation ratios $\alpha_{r}$ and $\alpha_{c}$ with a fixed $p=6$ and $N$ varying from 10 to 26. The performance with fixed parameters remains consistently good, even for larger problem sizes. The shaded regions represent the standard errors.
@@
This set of fixed angle is a good parameter for other problem instances from the family of $3$-regular graphs. We define the performance ratio as the approximation ratio of the QAOA-for-QRAO with fixed parameters over the one with optimized parameter. In Fig. 5A, we see that the performance ratio of both $\alpha_{r}$ and $\alpha_{c}$ are close to 1, especially with rounding the fixed angle performance could be better than the optimized parameter, indicting that it is sufficient to use a set of fixed angles for solving QAOA-for-QRAO. In Fig. 5B, we show $\alpha_{r}$ and $\alpha_{c}$ growing as $p$ with fixed parameters, confirming that QAOA-for-QRAO is able to converge to the ground state of $\tilde{H}_{C}$ as $p$ increases. Regarding the $\alpha_{c}$, generally it is positively correlated to $\alpha_{r}$ depending on different approaches of rounding. We refer the detailed discussion to Ref. [7, 8], and leave a better rounding approach, such as the tree rounding [18], as a future work. In Fig. 5C, we show that the fix angles work well for problems of size up to $N=26$ that is beyond the size of problems where we obtained the average parameters. In summary, this implies the parameter setting heuristics of standard QAOA still hold for QAOA-for-QRAO. Since we used no QRAO-specific structure, it is likely that same parameter setting heuristics will also work well for QAOA applied to problems in quantum chemistry and beyond.
@@ III-C Cost Hamiltonian Implementation
$e^{-i\gamma\tilde{H}_{C}}$ $=$ $\exp{-i\gamma\sum_{(i,j)\in E}P^{(q_{i})}_{i}P^{(q_{j})}_{j}}$ ) (13)
@@
$\approx$ ${\left[\prod_{(q_{i},q_{j})}\exp{-i\frac{\gamma}{T}\sum_{\begin{subarray}{c}(a,b)\in E\\ q_{a}=q_{i}\land q_{b}=q_{j}\end{subarray}}P^{(q_{i})}_{a}P^{(q_{j})}_{b}}\right]}^{T}$
@@ IV Comparison to the standard QAOA
In Fig. 7 A, we report the average results $\alpha_{c}$ of 10 random instances of 3-regular with $N=20$ nodes. In Fig. 7 B, we show the average $\alpha_{c}$ at $p=6$ under varying noise levels. The error probabilities in the emulator are simultaneously scaled by a constant factor. A noise scale of 1 corresponds to the noise level of the current H1-1 device emulator. As the noise scale increases, the simulation is conducted under noisier conditions. For the same graph instances, QAOA outperforms QAOA-for-QRAO in both noiseless and noisy scenarios. As shown in Fig. 5 and  6, the value of $\alpha_{c}$ can be smaller than $\alpha_{r}$ due to the rounding process. However, our goal is not to claim that QAOA-for-QRAO is superior to standard QAOA in terms of solution quality $\alpha_{c}$. Instead, we highlight the trade-off between quantum resource usage and algorithm performance. Notably, the QAOA-for-QRAO formulation requires up to three times fewer qubits than QAOA. We plot the number of qubits needed to solve the 3-regular graphs in Fig. 7 C. Future improvements in rounding techniques and more customized QAOA designs (e.g., Hamiltonian simulation and parameter fine-tuning) could potentially narrow the performance gap between QAOA-for-QRAO and QAOA

### v2

@@ Abstract
Solving hard optimization problems is one of the most promising application domains for quantum computers due to the ubiquity of such problems in industry and the availability of broadly applicable theoretical quantum speedups. However, the ability of near-term quantum computers to tackle industrial-scale optimization problems is limited by their size and the overheads of quantum error correction. Quantum Random Access Optimization (QRAO) has been proposed to reduce the space requirements of quantum optimization. However, to date QRAO has only been implemented using variational algorithms, which suffer from the need to train instance-specific variational parameters, making them difficult to scale. We propose and benchmark a non-variational approach to QRAO based on the Quantum Alternating Operator Ansatz (QAOA) for the MaxCut problem. We show that instance-independent “fixed” parameters achieve good performance, removing the need for variational parameter optimization. Additionally, we evaluate different design choices, such as various mixers, initial states, and QRAO-specific implementations of the QAOA cost operator, and identify a strategy that performs well in practice. Our results pave the way for the practical execution of QRAO on early fault-tolerant quantum computers.
@@ I Introduction
Combinatorial optimization (CO) is considered a promising application domain for quantum computers due to the ubiquity of complex optimization problems in industry and the potential for quantum algorithmic improvements [1, 2, 3, 4, 5, 6]. Recent resource estimates suggest that quantum approximate optimization algorithm may achieve a speedup over state-of-the-art classical solvers on realistic future fault-tolerant quantum computers [7]. However, achieving practical quantum advantage still faces numerous challenges, including the overhead associated with error correction and the necessity for more substantial algorithmic speedups. Among these, a significant challenge for near-term quantum computers is the limited number of computational qubits, stemming from difficulties in scaling quantum devices and the considerable space required for error correction.
Quantum Random Access Optimization (QRAO) [8, 9] is a quantum-relaxation based optimization method that reduces the space requirements of quantum optimization algorithms by encoding more than one variable per qubit. It utilizes the quantum superposition property, namely, a qubit can be in a superposition of 0 and 1 states, and turns the noise into advantage. That is, while a qubit is necessary to encode a binary variable without error by Holevo bound, it can encode more variables if error is tolerated. Quantum Random Access Codes (QRACs) [10, 11] are the family of such encodings; $(m,N,p)$-QRAC is an encoding of $m$ binary variables on $N$ qubits so that any 1 out of $m$ can be extracted with probability $p>1/2$ (we may skip $p$ in the QRAC notation for simplicity). It is known that there exist $(4^{N}-1,N,p)$-QRACs but not $(4^{N},N,p)$-QRACs [12], while optimal construction of QRACs is only known for $N=1,2$ [12, 13, 14]. QRACs were first used for CO in [8], extended theoretically in [9], and since then there have been various research revealing QRAO’s advantages; it can scale up the CO instances solvable on near-term quantum devices [15, 16], be robust to noise [17], applicable to constrained CO [18], and its recursive variant can achieve better solutions than other state-of-the-art quantum approaches [19].
@@ II-B Quantum Random Access Optimization (QRAO) formulation
QRAO is based on the idea of encoding multiple binary variables in a qubit. In this paper, we focus on encoding 3 binary variables into a qubit using the $(3,1,1/2+1/(2\sqrt{3}))$-QRAC that encodes the bit string $b_{1}b_{2}b_{3}\in\{0,1\}^{3}$ into the quantum state (in the density matrix representation [31])
@@
The main idea of QRAO for solving a $3N$-variable CO instance is to construct a cost Hamiltonian $-\tilde{H}_{C}$ of an $N$-qubit system so that its ground state closely resembles $\rho_{b^{*}}\equiv\rho_{b_{1}^{*}b_{2}^{*}b_{3}^{*}}\otimes\cdots\otimes\rho_{b_{3N-2}^{*}b_{3N-1}^{*}b_{3N}^{*}}$, where $b^{*}$ is the optimal solution of the CO instance. The cost Hamiltonian of QRAO for MaxCut $G(V,E)$, here denoted by $\tilde{H}_{C}$, is easily generalized from that of QAOA in Eq. (2) as
@@
However, because the ground state $\rho$ of $-\tilde{H}_{C}$ is likely to be a (pure) entangled state found on larger subspace than $\rho_{b}$, $\mbox{Tr}\left(\rho\tilde{H}_{C}\right)\geq\mbox{Tr}\left(\rho_{b}\tilde{H}_{C}\right)$ holds and some rounding procedure, such as the Pauli rounding, must be performed to find the closest $\rho_{b}$ to $\rho$. For this reason, $\tilde{H}_{C}$ is also called relaxed Hamiltonian. Other rounding methods are the magic rounding [8], that randomly projects each qubit of $\rho$ into one-out-of-eight state of Eq. (8), and the tree rounding [19].
@@ III-B Parameter Setting Protocol
Figure 5: Performance of QAOA-for-QRAO with fixed parameters. A We present the ratio of the performance with fixed parameters to the performance with optimized parameters. The fixed parameters performance is comparable to the ones with optimized parameters across different values of $p$ in both $\alpha_{c}$ (blue line) and $\alpha_{r}$ (orange line). B We show the average approximation ratios $\alpha_{r}$ and $\alpha_{c}$ for 120 random instances, with $N$ varying from 10 to 16 and $p$ up to 6. The performance improves as $p$ increases, indicating that the fixed parameters are of high quality across different values of $p$. C We present the average approximation ratios $\alpha_{r}$ and $\alpha_{c}$ with a fixed $p=6$ and $N$ varying from 10 to 26. The performance with fixed parameters remains consistently good, even for larger problem sizes. The shaded regions represent the standard errors.
@@
This set of fixed angle is a good parameter for other problem instances from the family of $3$-regular graphs. We define the performance ratio as the approximation ratio of the QAOA-for-QRAO with fixed parameters over the one with optimized parameter. In Fig. 5A, we see that the performance ratio of both $\alpha_{r}$ and $\alpha_{c}$ are close to 1, especially with rounding the fixed angle performance could be better than the optimized parameter, indicting that it is sufficient to use a set of fixed angles for solving QAOA-for-QRAO. In Fig. 5B, we show $\alpha_{r}$ and $\alpha_{c}$ growing as $p$ with fixed parameters, confirming that QAOA-for-QRAO is able to converge to the ground state of $-\tilde{H}_{C}$ as $p$ increases. Regarding the $\alpha_{c}$, generally it is positively correlated to $\alpha_{r}$ depending on different approaches of rounding. We refer the detailed discussion to Ref. [8, 9], and leave a better rounding approach, such as the tree rounding [19], as a future work. In Fig. 5C, we show that the fix angles work well for problems of size up to $N=26$ that is beyond the size of problems where we obtained the average parameters. We also report the performance of fixed parameters under the $(2,1)$-QRAC in appendix A, where the observations still hold. In summary, this shows evidence the parameter setting heuristics of standard QAOA still hold for QAOA-for-QRAO. Given that the parameter setting heuristics of QAOA have been successfully validated across various problems, including chemistry [27], the low-correlation binary sequence problem [3], the k-SAT problem [2], and the Sherrington-Kirkpatrick model [42], it is likely that these heuristics will perform well for QAOA-for-QRAO in other problems. We propose to explore the validation of these heuristics in additional problems in future research.
@@ III-C Cost Hamiltonian Implementation
$e^{-i\gamma\tilde{H}_{C}}=\exp{-i\gamma\sum_{(i,j)\in E}P^{(q_{i})}_{i}P^{(q_{j})}_{j}}$ )
@@
$\approx{\left[\prod_{(q_{i},q_{j})}\exp{-i\frac{\gamma}{T}\sum_{\begin{subarray}{c}(a,b)\in E\\ q_{a}=q_{i}\land q_{b}=q_{j}\end{subarray}}P^{(q_{i})}_{a}P^{(q_{j})}_{b}}\right]}^{T}$ (13)
@@ IV Comparison to the standard QAOA
In Fig. 7 A, we report the average results $\alpha_{c}$ of 10 random instances of 3-regular with $N=20$ nodes. In Fig. 7 B, we show the average $\alpha_{c}$ at $p=6$ under varying noise levels. The error probabilities in the emulator are simultaneously scaled by a constant factor. A noise scale of 1 corresponds to the noise level of the current H1-1 device emulator. As the noise scale increases, the simulation is conducted under noisier conditions. For the same graph instances, QAOA outperforms QAOA-for-QRAO in both noiseless and noisy scenarios. As our QAOA-for-QRAO ansatz could have more two-qubit gates than the standard QAOA due to the overhead overhead associated with the Trotter approximation of the cost Hamiltonian, the gap between QAOA-for-QRAO and the standard QAOA becomes larger the noise strength increases. As shown in Fig. 5 and  6, the value of $\alpha_{c}$ can be smaller than $\alpha_{r}$ due to the rounding process. However, our goal is not to claim that QAOA-for-QRAO is superior to standard QAOA in terms of solution quality $\alpha_{c}$. Instead, we highlight the trade-off between quantum resource usage and algorithm performance. Notably, the QAOA-for-QRAO formulation requires up to three times fewer qubits than QAOA. We plot the number of qubits needed to solve the 3-regular graphs in Fig. 7 C. Future improvements in rounding techniques and more customized QAOA designs (e.g., Hamiltonian simulation and parameter fine-tuning) could potentially narrow the performance gap between QAOA-for-QRAO and QAOA
[…]
@@ V Conclusion
Data availability
The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.
Code availability
The code for numerical simulations is available upon reasonable request.
Appendix A Performance with $(2,1)$-QRAC
We solve the same MaxCut problem instances with QAOA-for-QRAO using $(2,1)$-QRAC that encodes 2 binary variables into a qubit. We follow the same protocol to extract a fixed set of parameters and use Y mixer for the QAOA-for-QRAO ansatz. In Fig. 9, we report the results on 120 random instances with N varying from 10 to 16 and p up to 6. It shows that a fixed parameter still performs well and the observations are consistent with the ones from $(3,1)$-QRAC experiments.
Figure 9: Parameter schedule and performance of QAOA-for-QRAO with $(2,1)$-QRAC. The observations on $(3,1)$-QRAC still hold. Left: Optimized and fixed parameter schedules of one $N=16$ instance at $p=6$ with $(2,1)$-QRAC. Right: The ratio of the performance with fixed parameters to the performance with optimized parameters. The fixed parameters performance is comparable to the ones with optimized parameters across different values of $p$ in both $\alpha_{c}$ (blue line) and $\alpha_{r}$ (orange line).